OpenStudy (anonymous):

I need help! Which of the following is an equivalent form of the compound inequality −44 > −2x − 8 ≥ −8? −2x − 8 > −44 and −2x − 8 ≥ −8 −2x > −44 and −8 ≥ −8 −2x − 8 < −44 and −2x − 8 ≤ −8 −2x − 8 < −44 and −2x − 8 ≥ −8

OpenStudy (***[isuru]***):

hmmm... break the inequality in to 2 parts.. then solve then separately.. It will lead u to your answer :)

OpenStudy (***[isuru]***):

\[-44 > -2x -8 \ge -8 = -44 > -2x -8 \ and \ -2x -8 \ge -8\]